OpenStudy (anonymous):

Solve this DE by seperation of variables: yln(x)x'=(y+1)^2/x^2

OpenStudy (anonymous):

alright i split it up to get \[yln(x)\frac{dx}{dy}=\frac{(y+1)^2}{x^2}\]

OpenStudy (anonymous):

then made a common factor to fact out CF=\[\frac{x^2}{y}dy\] which got me to \[x^2ln(x)dx=\frac{(y+1)^2}{y}dy\]

OpenStudy (anonymous):

squaring the top and seperating the numerator while also taking anti derivative \[\int\limits_{}^{}x^2\ln(x)dx=\int\limits_{}^{}\frac{y^2}{y}dy+\int\limits_{}^{}\frac{2y}{y}dy+\int\limits_{}^{}\frac{dy}{y}\]

OpenStudy (anonymous):

simplifying and doing the right side you get \[\int\limits_{}^{}x^2\ln(x)dx=\frac{y^2}{2}+2y+\ln(y)+c\]

OpenStudy (anonymous):

now doing the right side by parts is where i got something different from the book

OpenStudy (anonymous):

using by parts i get \[u=ln(x)\] \[du=\frac{1}{x}\]

OpenStudy (anonymous):

\[dv=x^2\] \[v=\frac{x^3}{3}\]

OpenStudy (anonymous):

\[uv-\int\limits_{}^{}vdu=\frac{x^3}{3}\ln(x)-\int\limits_{}{}\frac{x^2}{x}dx\]

OpenStudy (anonymous):

]

OpenStudy (anonymous):

not sure what happened there sam as i can't see anything

OpenStudy (anonymous):

simplifying the integral and doing it yields \[\frac{x^3}{3}\ln(x)-\frac{x^2}{2}+c_2\]

OpenStudy (anonymous):

but that is not what is within the book?

OpenStudy (anonymous):

trying to notice my mistake

OpenStudy (anonymous):

or if it's a mistake in the book

OpenStudy (anonymous):

the y side is correct however this is what they have for the answer \[\frac{x^3}{3}\ln(x)-\frac{x^3}{9}=\frac{y^2}{2}+2y+\ln(y)+c\]

OpenStudy (ash2326):

Sorry I thought you are doing it by using integrating factor:(

OpenStudy (anonymous):

i found the error instead of x^3 i had x^2 only idk why but thanks anyway

OpenStudy (anonymous):

somewhere inbetween i must have mixed it up

OpenStudy (anonymous):

\[\int\limits_{}{}\frac{x^3}{3x}dx=\frac{1}{3}\int\limits_{}{}x^2dx=\frac{1}{9}x^3\]

OpenStudy (ash2326):

Seems right:)

OpenStudy (anonymous):

\[\frac{1}{3}x^3(\ln(x)-\frac{1}{3})\]